If at one atmospheric pressure the gas is spreading from 20 cm cube to 50 cm cube at constant temperature then find the final pressure

Grade:11

2 Answers

Arun
5 years ago
Dear student
 
Boyle’s law
 
P1 * V1 = P2 * V2
 
1 * 20 = P2 * 50
 
P2 = 0.4 atm
 
Regards
Arun (askIITians forum expert)
Khimraj
5 years ago
As the temperature is constant
sofrom ideal gas equation
PV = constant
then
Pi*Vi = Pf*Vf
1*20 = Pf*50
so 
Pf = 0.4 atm
Hope it clears.